Status Report

NASA Space Station On-Orbit Status 4 September 2004

By SpaceRef Editor
September 4, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 4 September 2004

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.  Saturday, with the crew awaking to kudos for their very smooth 5h 21m spacewalk yesterday.

After their long day yesterday, Padalka and Fincke slept in today, with wakeup slipped by 7 hrs to 9:00am EDT, followed by a short 8.5-hr day.  Sleep time is 5:30pm tonight.

All ISS systems are back in a normal configuration with post-EVA reconfiguration complete, including comm links, OpsLAN network, IMV (inter-modular ventilation), TCCS (trace contaminant control subsystem), ITCS (internal thermal control system), and PCUs (plasma contactor units) back to Standby. 

Also, the station is once again flying in sun-pointing XPOP (x-axis perpendicular to orbit plane) attitude.

Regular post-EVA activities by both crewmembers this morning included removing BK3 oxygen tanks and batteries from their Orlan suits’ backpacks, refilling the spacesuits’ feedwater bladders with water, and setting the Orlans up for drying out.

In addition, CDR Padalka reactivated the new Sputnik-SM Kenwood D700 amateur radio station in the Service Module (SM) and stowed the “Urolux” equipment used by the crew last night for the obligatory post-EVA session of the Russian crew health-monitoring program’s MO-9/Biochemical Urinalysis.

Gennady also returned the Kriogem-03M refrigerator equipment from the SM Transfer Tunnel (PrK) to the DC1 docking compartment.   [The cooler currently contains the Japanese JCF-01V crystallizer unit of the GCF (Granada Crystallization Facility) and LUCH-2 biological crystallization hardware.]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

FE/SO Fincke brought the SODF (Systems Operation Data File) and Emergency books plus CD Library from their temporary location in the DC1 docking module back to their regular place in the Lab, then demated and removed the UOP DCP (utility outlet panel/display & control panel) bypass power cable at the Lab RWS (robotics work station) that was used to support video camera coverage of the EVA.

Mike completed the regular once-a-week maintenance reboot on the operational PCS laptops and the restart of the OCA comm router laptop (every two weeks).

The FE also conducted today’s routine inspection of the SM’s SOZh life support system (including replacement of ASU toilet facility inserts).

For his “Saturday Science” program this weekend, the Science Officer had selected the HEAT experiment, a two-day operation.  Today, Mike tagged up with the ESA Principal Investigator of HEAT and an MSG (Microgravity Science Glovebox) specialist at POIC (Payload Operations & Integration Center, MSFC) to review tomorrow’s experiment session.   [The European HEAT was one of three experiments of Andre Kuipers’ VC-6 Delta science program that were performed in the USOS.  Installed in the MSG, it was to investigate heat transfer performance of grooved heat pipes, but the experiment failed to work properly during the VC-6 joint crew period, hampered by insufficient heat transfer from the condenser area of the HEAT base plate to the MSG coldplate, upon which HEAT is mounted.  Mike has now volunteered to run it again.  The experiment’s main scientific objective is to characterize the heat transfer performance of a grooved heat pipe under micro-G conditions by deriving the maximum sustainable heat flux in various heating and cooling modes.  Investigators hope to prove that the grooved heat pipe design is more appropriate than a sleek design.]

Padalka completed another daily inspection of the BIO-5 Rasteniya-2 (“Plants-2”) plant growth experiment in the Lada-5 greenhouse.  Topping its water container off as required was an additional step included in the crew-choice task list.

Planning has started on trash gathering for Soyuz 8S and Progress 15P.   [Until it is determined which items will go in which vehicle, the crew will gather and pack approved items in bags, with their destinations to be decided later by TsUP.  Some trash packing activities will be included in the standard housekeeping time each week.]

Objects jettisoned yesterday during EVA-11 (retrograde, i.e., slowing down and dropping in altitude):

  • Connector cap (from FGB zenith)
  • WAL-1 antenna cover (from SM aft)
  • WAS antenna cover (from SM aft)
  • WAS antenna cover (from SM aft)
  • WAL-3 antenna cover (from SM aft)
  • WAL-2 antenna cover (from SM aft)
  • Cleaning towels (from SM stbd, near foot restraint)
  • PIG pressurized container with RRZh valve panel (from DC1 stbd aft)
  • possibly three more connector caps.

More information and background material on EVA 4 can be found at the Expedition 9 EVA 4 Reference Page

Weekly Science Update (Expedition Nine — 19th):

GASMAP:  Nothing new.

Human Research Facility/Workstation (HRF WS):  Continuing.

Advanced Ultrasound (ADUM):  Nothing new.

Hand Posture Analyzer (HPA:  Nothing new.

In-Space Soldering Investigation (ISSE):  Nothing new.

Pore Formation and Mobility Investigation (PFMI):  Nothing new.

Space Acceleration Measurement System (SAMS):  Nothing new.

Microgravity Acceleration Measurement System (MAMS):  Nothing new.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):  Nominal.

Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS):  Nothing new.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):    Planned.

Binary Colloidal Alloy Test-3 (BCAT-3):   Complete.

Renal Stone (RS):  Nothing new.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES):  Nothing new.

Foot/Ground Reaction Forces during Space Flight (FOOT):  Nothing new.

Materials ISS Experiment (MISSE):  In progress. Deployed outside on the U.S. Airlock.  Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI):  Nothing new.

Biotechnology Specimen Temperature Controller (BSTC):  In planning.

Yeast Group Activation Packs (Yeast GAP):  Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM):  Successful camera battery charge and function test on 8/28. Thanks!

Earth Science Toward Exploration Research (ESTER):  Nothing new.

Serial Network Flow Monitor (SNFM):  Nothing new.

Fluid Merging Viscosity Measurement (FMVM):  Nothing new.

Viscous Liquid Foam–Bulk Metallic Glass (Foam):   Nothing new.

BIOPSY (Effects of Prolonged Space Flight on Human Skeletal Muscle):  Nothing new.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2):  Planned.

Commercial Generic Bioprocessing Apparatus (CGBA):  Nothing new.

Miscible Fluids in Microgravity (MFMG):  Nothing new.

Educational Payload Operations (EPO):  EPO looks forward to future operations.  The crew’s previous EPO demonstrations have been well received by NASA Education and payload sponsors.

Capillary Flow Experiment (CFE):    Nothing new.

Crew Earth Observations (CEO):  A striking ISS/CEO image of the Ksudach Volcano on the Kamchatka Peninsula of far eastern Russia will be featured on NASA’s Earth Observatory website this weekend.  Such detailed views of this area are rare because of poor weather and lighting conditions.  Images of Hurricane Frances have dominated this week’s targets and imagery reviews.  With the crew’s great response and prompt downlinks, ground personnel has been able post their images on their website for the world to see in less than 10 hours after downlink.  PAO, the National Hurricane Center, and the news media have all been pleased to have timely access to these images.  Wednesday morning’s sun glint views are the most beautiful and dramatic seen so far from the ISS.  Investigators are continuing to review and shorten their standing target list for this increment as the crew’s success in acquiring them continues.  Expedition 9 image total is now approaching 15,000.

Today’s optional CEO photo targets, limited in the current XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in ram), were Hurricane Howard, Baja California (Dynamic Event.  Howard is now a Category 3 storm and should display a well-developed eye and banding.  Looking approximately 3 deg to the left of track.  The storm is not currently predicted to make landfall), Internal waves, Tuamotu Archipelago (waters to the NE of the archipelago should have been visible for internal wave photography. The sunglint point was left of track), and Typhoon Songda, south of Japan (Dynamic Event.  This typhoon is bearing west towards Taiwan.  The storm is larger than Hurricane Frances.  It is large enough that the crew should be able to take some dramatic panoramas.  Looking approximately 10 deg right of track).

CEO images can be viewed at these websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 9 crew visit:

Major upcoming events:

  • Reboost — 9/22 (phase angle correction for 9S)
  • Soyuz 9S launch — 10/9;
  • Soyuz 9S dock — 10/11;
  • Soyuz 8S undock/land — 10/19;
  • Soyuz 9S relocate — 11/18;
  • Progress 16P launch — 11/24.

Expedition 9 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this noon, 1:35pm EDT [= epoch]):

  • Mean altitude — 360.7 km
  • Apogee height — 364.6 km
  • Perigee height — 356.8 km
  • Period — 91.76 min.
  • Inclination (to Equator) — 51.63 deg
  • Eccentricity — 0.0005766
  • Solar Beta Angle — 49.3 deg (magnitude increasing)
  • Orbits per 24-hr. day — 15.69
  • Mean altitude loss in last 24 hours — 130 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 33081


ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.